Question 1 (1 point) (1 Point) We assume that the representative consumers preferences exhibit the properties that Oa) theyQuestion 9 (1 point) (1 Point) As the quantity of capital increases, the marginal product of labour Oa) is constant. Ob) may
Answer #1

Q1. Option c

As quantity demanded decreases with the increase in price and consumer has varied tastes, preferences

Q2. Option b

As consumer substitute with superior good with the increase in income

Q3. Option b

It is the opportunity cost of gaining an additional quantity of a good by forgoing another additional quantity of another good

Q4. Option a

All agents spend whatever they have because there is no future to save

Q9. Option d

As increase in capital increases the productivity

Q10. Option c

Q11. Option b

Profit = TR-TC= 1200-(700+100+200) = 200

Q12. Option d

As it may discourage people from consumption